OpenStudy (becky):

find the points p and q find the distance d(pq) the corrdinates of the mid-point of the segment po 1. P(-4,3), Q(2,-6)

OpenStudy (mathmate):

Hints: mid-point formula: \(M(x,y)=M(\frac{x1+y1}{2},\frac{y1+y2}{2})\) Distance formula: \(D(PQ)=\sqrt{(x2-x1)^2+(y2-y1)^2}\)